OpenStudy (anonymous):

for the following equation y=-1/2x+5 complete the following ordered pair so that it is a solution to the equation ( , -3

OpenStudy (anonymous):

You have a y value of -3 so plug that in for y and solve for x -3 = -1/2x + 5

OpenStudy (anonymous):

Subtract 5 from both sides -5 = -1/2x

OpenStudy (anonymous):

Multiply both sides by -2 10 = x

OpenStudy (anonymous):

so (10,-3)

OpenStudy (anonymous):

isolate x -1/2 x = y -5 x=(y-5)/-1/2 repalace y=-3 into this equation x=16 so the point is (16, -3)